MINDTAP Search this course From Cengage Chapter 2 Assignment Check My Work 2. 3 4 3. eBook Problem 2-23 approximately 100. Answer the following questions using a bell-shaped distribution and the empirical rule for the math test scores If required, round your answers to two decimal places. (a) What percentage of students have an SAT math score greater than 621? o Suppose that the national average for the math portion of the College Boards SAT is 521The Cllege Board periodically rescales the test scores such that the standard deviation is (b) What percentage of students have an SAT math score greater than 721? (c) What percentage of students have an SAT math score between 421 and 521? (d) What is the z-score for student with an SAT math score of 635? (e) What is the z-score for a student with an SAT math score of 425? Check My Work 0-Icon Kgy
Answer #1

a)as 621 is 1 std deviation above mean ; therefore % of score above 621 =16%

b)as 721 is 2 std deviation above mean ; therefore % of score above 721 =2.5%

c)% of score between 421 and 521 =34.0%

d)z score =(X-mean)/std deviation=(635-521)/100=1.14

e)z score=(425-521)/100=-0.96